# Inflationary Models with Gauss–Bonnet Coupling in Light of ACT Observations

### v1

@@ 3 The model
Given a potential $V(\phi)$, choosing the Gauss–Bonnet coupling function as [52]
$\xi(\phi)=\frac{3\lambda}{4V(\phi)},$ (3.1)
then, to the first order of the slow-roll parameters, the slow-roll parameters satisfy the relations
$\delta_{1}=2\lambda\epsilon_{1},\quad\delta_{i+1}=\epsilon_{i+1},\quad i\geq 1,$ (3.2)
and the expression for $Q$ in Eq. (2.15) reduces to
$Q=(1-\lambda)\frac{V_{,\phi}}{V}.$ (3.3)
Using Eq. (3.2) and the expressions in Eq. (2.12), the scalar spectral index and the tensor-to-scalar ratio become
$n_{s}-1=(1-\lambda)\left(2\eta_{V}-6\epsilon_{V}\right),$ (3.4)
$r=16(1-\lambda)^{2}\epsilon_{V},$ (3.5)
where $\epsilon_{V}$ and $\eta_{V}$ are the standard potential slow-roll parameters in the canonical inflation model,
$\epsilon_{V}=\frac{1}{2}\left(\frac{V_{,\phi}}{V}\right)^{2}~,~\eta_{V}=\frac{V_{,\phi\phi}}{V}.$ (3.6)
To the first order of the slow-roll parameters, the relation of $e$-folding number $N$ given by Eq. (2.16) becomes
$N=\frac{1}{1-\lambda}\int_{\phi_{e}}^{\phi}\frac{V}{V_{,\phi}}d\phi.$ (3.7)
@@
$\phi=\sqrt{1-\lambda}\varphi,$ (3.8)
@@
$U(\varphi)=V(\phi)=V[\phi(\varphi)].$ (3.9)
In terms of the new field $\varphi$ and potential $U(\varphi)$, to the first order of the slow-roll parameters, the scalar spectral index (3.4) and tensor-to-scalar ratio (3.5) reduce to
$n_{s}-1=\left(2\eta_{U}-6\epsilon_{U}\right),$ (3.10)
$r=16(1-\lambda)\epsilon_{U},$ (3.11)
@@
$\epsilon_{U}=\frac{1}{2}\left(\frac{U_{,\varphi}}{U}\right)^{2}~,~\eta_{U}=\frac{U_{,\varphi\varphi}}{U},$ (3.12)
and the $e$-folding number $N$ given in Eq. (3.7) takes the form
$N=\int_{\varphi_{e}}^{\varphi}\frac{U}{U_{,\varphi}}d\varphi.$ (3.13)
Therefore, with the Gauss-Bonnet coupling function of the form $\xi(\phi)=3\lambda/[4V(\phi)]$, the scalar spectral index remains identical to that from the corresponding canonical model with the same potential and a rescaled field, while the tensor-to-scalar ratio is suppressed by a factor of $(1-\lambda)$. As a result, $r$ can be made sufficiently small without affecting $n_{s}$. This effectively removes the observational constraint on the tensor-to-scalar ratio, allowing model comparisons with data to focus solely on the scalar spectral index. In this way, one degree of observational tension is eliminated, making it significantly easier for inflationary models to comply with current data, such as the recent constraints from P-ACT-LB data [11, 12]:
$n_{s}=0.9743\pm 0.0034.$ (3.14)
@@ 3.1 Chaotic inflation
For chaotic inflation with a monomial potential of the form [61]
$V(\phi)=V_{0}\phi^{p},$ (3.15)
@@
$n_{s}=1-\frac{p+2}{2(N+\tilde{n})},\quad r=\frac{4(1-\lambda)p}{N+\tilde{n}}$, (3.17)
with $\tilde{n}=p/4$ for $2/3<p<2$, and $\tilde{n}=|p-1|/2$ for other cases. Taking $N=60$, to be consistent with the P-ACT-LB constraint on $n_{s}$, given in Eq. (3.14), the index of the chaotic inflation should satisfy
$0.68<p<1.51.$ (3.18)
The comparison between the model predictions for different power indices $p$, as given in Eq. (3.17), and the observational constraints from P-ACT-LB-BK18 is shown in Fig. 1. Here, P-ACT-LB-BK18 refers to the joint constraints obtained from the P-ACT-LB dataset combined with the B-mode polarization measurements from the BICEP/Keck Array experiments at the South Pole (BK18) [62]. The $1\sigma$ and $2\sigma$ confidence regions from the P-ACT-LB-BK18 data are indicated by the purple contours. The dashed curves represent predictions from inflationary models without Gauss-Bonnet coupling (i.e., $\lambda=0$), while the solid curves correspond to models with Gauss-Bonnet coupling with $\lambda=0.8$. In all cases, the number of $e$-folds is fixed to $N=60$. The black curves denote the predictions from the chaotic inflation model. Without the Gauss-Bonnet term, chaotic inflation is disfavored by the observational data. However, with Gauss-Bonnet coupling, the tensor-to-scalar ratio is significantly reduced, bringing the model predictions into agreement with current constraints.
@@ 3.2 $\alpha$ attractors
For the $\alpha$-attractor E-model with a potential of the form [7, 63]
$V(\phi)=V_{0}\left[1-\exp\left(-\sqrt{\frac{2}{3\alpha}}\phi\right)\right]^{2n},$ (3.19)
@@
$U(\varphi)=V_{0}\left[1-\exp\left(-\sqrt{\frac{2}{3\tilde{\alpha}}}\varphi\right)\right]^{2n},$ (3.20)
@@
$\tilde{\alpha}=\frac{\alpha}{1-\lambda}.$ (3.21)
The resulting expressions for the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ are [64]
$n_{s}=1+\frac{8n}{3\tilde{\alpha}\left[g(N,n,\tilde{\alpha})+1\right]}-\frac{8n(n+1)}{3\tilde{\alpha}\left[g(N,n,\tilde{\alpha})+1\right]^{2}},$ (3.22)
$r=\frac{64(1-\lambda)n^{2}}{3\tilde{\alpha}\left[g(N,n,\tilde{\alpha})+1\right]^{2}},$ (3.23)
@@
$g(N,n,\tilde{\alpha})=W_{-1}\left[-\left(\frac{2n}{\sqrt{3\tilde{\alpha}}}+1\right)\exp\left(\frac{-4nN}{3\tilde{\alpha}}-\frac{2n}{\sqrt{3\tilde{\alpha}}}-1\right)\right],$ (3.24)
@@
$n_{s}=1-\frac{2}{N},\quad r=\frac{12(1-\lambda)\tilde{\alpha}}{N^{2}},$ (3.27)
which are disfavored by the P-ACT-LB data. In the large-$\tilde{\alpha}$ limit, the E-model asymptotically approaches chaotic inflation with a power-law potential of index $p=2n$, and the predictions reduce to those given in Eq. (3.17). Taking $N=60$ and using Eqs. (3.14) and (3.22), the observational constraints on the parameters $\tilde{\alpha}$ and $n$ from the P-ACT-LB data are shown in Fig. 2. For $n=1/2$, the predictions of the E-model for different values of $\tilde{\alpha}$ are shown in Fig. 1, represented by the blue curves. With the inclusion of the Gauss-Bonnet term and a coupling constant $\lambda=0.8$, the E-model becomes consistent with the P-ACT-LB-BK18 observational data for sufficiently large $\tilde{\alpha}$.
Figure 2: Constraints on the parameters $n$ and $\tilde{\alpha}$ of the E-model, derived from the P-ACT-LB data as given in Eq. (3.14).
For the $\alpha$-attractor T-model with the potential [6, 7]
$V(\phi)=V_{0}\tanh^{2n}\left(\frac{\phi}{\sqrt{6\alpha}}\right),$ (3.28)
@@
with the effective parameter $\tilde{\alpha}$ also given by Eq. (3.21).
The predictions for the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ depend on the parameter regime. For the case with $n>1$ and $(4n^{2}-2n\sqrt{4n^{2}-1})/3<\tilde{\alpha}<{4n^{2}}/[3(n^{2}-1)]$, $1/\sqrt{3}<n<1$ and $\tilde{\alpha}>(4n^{2}-2n\sqrt{4n^{2}-1})/3$ or $1/3<n<1/\sqrt{3}$ and $\tilde{\alpha}>4n^{2}/[3(9n^{2}-1)]$, the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ are given by [65, 64]
$n_{s}=1-\frac{2}{N}+\frac{2N\sqrt{12n^{2}/\tilde{\alpha}+9}-6n(N-1)}{N\left[2N\sqrt{12n^{2}/\tilde{\alpha}+9}+n\left(4N^{2}/\tilde{\alpha}+3\right)\right]},$ (3.30)
$r=\frac{48(1-\lambda)n}{2N\sqrt{12n^{2}/\tilde{\alpha}+9}+n\left(4N^{2}/\tilde{\alpha}+3\right)}.$ (3.31)
For the case with $n>1$ and $\tilde{\alpha}>{4n^{2}}/[3(n^{2}-1)]$, the expressions become [65, 64]
$n_{s}=1-\frac{2}{N}+\frac{8n\left[(N-1)\sqrt{9\tilde{\alpha}^{2}+24\tilde{\alpha}n^{2}+4n^{2}}+6\tilde{\alpha}n-(3\tilde{\alpha}+2)nN+2n\right]}{N\left[\left(\sqrt{9\tilde{\alpha}^{2}+4(6\tilde{\alpha}+1)n^{2}}+4nN-2n\right)^{2}-9\tilde{\alpha}^{2}\right]},$ (3.32)
$r=\frac{192(1-\lambda)\tilde{\alpha}n^{2}}{\left(\sqrt{9\tilde{\alpha}^{2}+4(6\tilde{\alpha}+1)n^{2}}-2n+4nN\right)^{2}-9\tilde{\alpha}^{2}}.$ (3.33)
For all other parameter choices, the predictions are given by [65, 64]
$n_{s}=1-\frac{2}{N}+\frac{8n\left[(N+1)\sqrt{9\tilde{\alpha}^{2}+(4-24\tilde{\alpha})n^{2}}-6n\tilde{\alpha}-(3\tilde{\alpha}-2)nN+2n\right]}{N\left[\left(\sqrt{9\tilde{\alpha}^{2}+(4-24\tilde{\alpha})n^{2}}+4nN+2n\right)^{2}-9\tilde{\alpha}^{2}\right]},$ (3.34)
$r=\frac{192(1-\lambda)\tilde{\alpha}n^{2}}{\left[\sqrt{9\tilde{\alpha}^{2}+(4-24\tilde{\alpha})n^{2}}+n(4N+2)\right]^{2}-9\tilde{\alpha}^{2}}.$ (3.35)
Similar to the E-model, in the small-$\tilde{\alpha}$ limit, the predictions of the T-model reduce to those of the $\alpha$-attractor form given in Eq. (3.27). In the large-$\tilde{\alpha}$ limit, the T-model approaches the chaotic inflation model with a power-law potential of index $p=2n$ [65], and the predictions coincide with those in Eq. (3.17). Taking $N=60$, the observational constraints on the parameters $\tilde{\alpha}$ and $n$ of the T-model from the P-ACT-LB data are shown in Fig. 3. For $n=1/2$, the predictions of the T-model for different values of $\tilde{\alpha}$ are shown in Fig. 1, represented by the red curves. With the inclusion of the Gauss-Bonnet term and a coupling constant $\lambda=0.8$, the T-model becomes consistent with the P-ACT-LB-BK18 observational data for sufficiently large $\tilde{\alpha}$.
Figure 3: Constraints on the parameters $n$ and $\tilde{\alpha}$ of the T-model, derived from the P-ACT-LB data as given in Eq. (3.14)

### v2

@@ 3 The model
In general situations, the slow-roll parameters $\epsilon_{i}$ and $\delta_{i}$ are independent. In this paper, we set the condition
$\delta_{1}=2\lambda\epsilon_{1},$ (30)
with $0<\lambda<1$. Substituting this condition into the definitions of the slow-roll parameters (6) and (7), we have
$\delta_{i+1}=\epsilon_{i+1},\quad i\geq 1.$ (31)
Using Eqs. (30) and (31), the scalar spectral index (24) and tensor-to-scalar ratio (28) become
$n_{s}-1=-2\epsilon_{1}-\epsilon_{2},$ (32)
$r=16(1-\lambda)\epsilon_{1}.$ (33)
Thus the scalar spectral index coincides with that of the canonical case, while the tensor-to-scalar ratio $r$ is suppressed by a factor $1-\lambda$. In other words, under the relation (30), the Gauss–Bonnet term suppresses $r$ without affecting $n_{s}$, in terms of the Hubble flow slow-roll parameters.
Using the relations (12) and (13), the condition (30) is equivalent to Yi:2018gse
$\xi(\phi)=\frac{3\lambda}{4V(\phi)}.$ (34)
Higher order curvature corrections, in particular the Gauss–Bonnet coupling, naturally arise in string theory. In the low-energy effective action of the heterotic string theory, the Gauss–Bonnet term typically couples to the dilaton field with an exponential function Callan:1985ia ; Gross:1986mw ; Gasperini:1996fu . For an exponential potential, the string-motivated Gauss–Bonnet coupling takes the form given in Eq. (34), which leads to power-law inflation Guo:2009uk ; Jiang:2013gza . Furthermore, the coupling (34) implies the condition (30), resulting the reduction to the tensor-to-scalar ratio.
Since the Hubble slow-roll parameters are not directly expressed in terms of the potential, in the following, we rewrite the results using the potential slow-roll parameters. In terms of the standard potential slow-roll parameters in the canonical inflation model,
$\epsilon_{V}=\frac{1}{2}\left(\frac{V_{,\phi}}{V}\right)^{2}~,~\eta_{V}=\frac{V_{,\phi\phi}}{V},$ (35)
the scalar spectral index and the tensor-to-scalar ratio become
$n_{s}-1=(1-\lambda)\left(2\eta_{V}-6\epsilon_{V}\right),$ (36)
$r=16(1-\lambda)^{2}\epsilon_{V},$ (37)
and the expression for $Q$ in Eq. (14) reduces to
$Q=(1-\lambda)\frac{V_{,\phi}}{V}.$ (38)
To the first order of the slow-roll parameters, the relation of $e$-folding number $N$ given by Eq. (15) becomes
$N=\frac{1}{1-\lambda}\int_{\phi_{e}}^{\phi}\frac{V}{V_{,\phi}}d\phi.$ (39)
@@
$\phi=\sqrt{1-\lambda}\varphi,$ (40)
@@
$U(\varphi)=V(\phi)=V[\phi(\varphi)].$ (41)
In terms of the new field $\varphi$ and potential $U(\varphi)$, to the first order of the slow-roll parameters, the scalar spectral index (36) and tensor-to-scalar ratio (37) reduce to
$n_{s}-1=\left(2\eta_{U}-6\epsilon_{U}\right),$ (42)
$r=16(1-\lambda)\epsilon_{U},$ (43)
@@
$\epsilon_{U}=\frac{1}{2}\left(\frac{U_{,\varphi}}{U}\right)^{2},\quad\eta_{U}=\frac{U_{,\varphi\varphi}}{U},$ (44)
and the $e$-folding number $N$ given in Eq. (39) takes the form
$N=\int_{\varphi_{e}}^{\varphi}\frac{U}{U_{,\varphi}}d\varphi.$ (45)
Therefore, with the Gauss-Bonnet coupling function of the form $\xi(\phi)=3\lambda/[4V(\phi)]$, the scalar spectral index remains identical to that from the corresponding canonical model with the same potential and a rescaled field, while the tensor-to-scalar ratio is suppressed by a factor of $(1-\lambda)$. As a result, $r$ can be made sufficiently small without affecting $n_{s}$. This effectively removes the observational constraint on the tensor-to-scalar ratio, allowing model comparisons with data to focus solely on the scalar spectral index. In this way, one degree of observational tension is eliminated, making it significantly easier for inflationary models to comply with current data, such as the recent constraints from P-ACT-LB data ACT:2025fju ; ACT:2025tim :
$n_{s}=0.9743\pm 0.0034.$ (46)
It should be noted that, in the extreme limit $\lambda\to 1$, the model may suffer from fine-tuning issues. Moreover, since our results are derived under the first-order slow-roll approximation, if $1-\lambda$ becomes smaller than the first-order slow-roll parameters, the expressions for the scalar spectral index $n_{s}$, Eq. (42), and the tensor-to-scalar ratio $r$, Eq. (43), are no longer applicable, and higher-order corrections should be taken into account. However, in many cases, the magnitude of $\lambda$ does not need to be close to 1; reducing the tensor-to-scalar ratio by a factor of 2 to 3 is sufficient to match observations, i.e., $\lambda\sim 2/3$. In practical scenarios, this eliminates the two aforementioned issues.
@@ 3.1 Chaotic inflation
For chaotic inflation with a monomial potential of the form Linde:1983gd
$V(\phi)=V_{0}\phi^{p},$ (47)
@@
$n_{s}=1-\frac{p+2}{2(N+\tilde{n})},\quad r=\frac{4(1-\lambda)p}{N+\tilde{n}}$, (49)
with $\tilde{n}=p/4$ for $2/3<p<2$, and $\tilde{n}=|p-1|/2$ for other cases. Taking $N=60$, to be consistent with the P-ACT-LB constraint on $n_{s}$, given in Eq. (46), the index of the chaotic inflation should satisfy
$0.68<p<1.51.$ (50)
To be consistent with the joint constraints on $n_{s}$ and $r$ obtained from the P-ACT-LB dataset combined with the B-mode polarization measurements from the Background Imaging of Cosmic Extragalactic Polarization (BICEP)/Keck Array experiments (BK18) BICEP:2021xfz , denoted as P-ACT-LB-BK18, we show in Fig. 1 the allowed parameter space of chaotic inflation with the Gauss–Bonnet coupling for $N=60$. The parameter values in the blue and gray regions predict $n_{s}$ and $r$ consistent with the $1\sigma$ and $2\sigma$ confidence regions from the P-ACT-LB-BK18 data, respectively, indicated by the purple contours, as shown in Fig. 2. Compared with using only the scalar spectral index $n_{s}$, the joint constraint with $r$ broadens the allowed range of the power index $p$.
Figure 1: Constraints on chaotic inflation with a Gauss–Bonnet coupling for $N=60$. The blue and gray regions correspond to parameter values predicting $n_{s}$ and $r$ consistent with the $1\sigma$ and $2\sigma$ confidence regions from the P-ACT-LB-BK18 data, respectively. The P-ACT-LB-BK18 data are shown in Fig. 2.
The direct comparison between the model predictions for different power indices $p$, as given in Eq. (49), and the observational constraints from P-ACT-LB-BK18 is shown in Fig. 2. The dashed curves represent predictions from inflationary models without Gauss-Bonnet coupling (i.e., $\lambda=0$), while the solid curves correspond to models with Gauss-Bonnet coupling with $\lambda=0.8$. In all cases, the number of $e$-folds is fixed to $N=60$. The black curves denote the predictions from the chaotic inflation model. Without the Gauss-Bonnet term, chaotic inflation is disfavored by the observational data. However, with Gauss-Bonnet coupling, the tensor-to-scalar ratio is significantly reduced, bringing the model predictions into agreement with current constraints.
@@ 3.2 $\alpha$ attractors
For the $\alpha$-attractor E-model with a potential of the form Kallosh:2013maa ; Carrasco:2015rva
$V(\phi)=V_{0}\left[1-\exp\left(-\sqrt{\frac{2}{3\alpha}}\phi\right)\right]^{2n},$ (51)
@@
$U(\varphi)=V_{0}\left[1-\exp\left(-\sqrt{\frac{2}{3\tilde{\alpha}}}\varphi\right)\right]^{2n},$ (52)
@@
$\tilde{\alpha}=\frac{\alpha}{1-\lambda}.$ (53)
The resulting expressions for the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ are Yi:2016jqr
$n_{s}=1+\frac{8n}{3\tilde{\alpha}\left[g(N,n,\tilde{\alpha})+1\right]}-\frac{8n(n+1)}{3\tilde{\alpha}\left[g(N,n,\tilde{\alpha})+1\right]^{2}},$ (54)
$r=\frac{64(1-\lambda)n^{2}}{3\tilde{\alpha}\left[g(N,n,\tilde{\alpha})+1\right]^{2}},$ (55)
@@
[…]
$\times\exp\left(\frac{-4nN}{3\tilde{\alpha}}-\frac{2n}{\sqrt{3\tilde{\alpha}}}-1\right)\bigg],$
@@
$n_{s}=1-\frac{2}{N},\quad r=\frac{12(1-\lambda)\tilde{\alpha}}{N^{2}},$ (59)
which are disfavored by the P-ACT-LB data. In the large-$\tilde{\alpha}$ limit, the E-model asymptotically approaches chaotic inflation with a power-law potential of index $p=2n$, and the predictions reduce to those given in Eq. (49). Taking $N=60$ and using Eqs. (46) and (54), the observational constraints on the parameters $\tilde{\alpha}$ and $n$ from the P-ACT-LB data are shown in Fig. 3. For $n=1/2$, the predictions of the E-model for different values of $\tilde{\alpha}$ are shown in Fig. 2, represented by the blue curves. With the inclusion of the Gauss-Bonnet term and a coupling constant $\lambda=0.8$, the E-model becomes consistent with the P-ACT-LB-BK18 observational data for sufficiently large $\tilde{\alpha}$.
Figure 3: Constraints on the parameters $n$ and $\tilde{\alpha}$ of the E-model, derived from the P-ACT-LB data as given in Eq. (46). The blue region corresponds to parameter values for which the predicted scalar spectral index given by Eq. (54) from the E-model is consistent with the P-ACT-LB data.
For the $\alpha$-attractor T-model with the potential Kallosh:2013hoa ; Kallosh:2013maa
$V(\phi)=V_{0}\tanh^{2n}\left(\frac{\phi}{\sqrt{6\alpha}}\right),$ (60)
@@
with the effective parameter $\tilde{\alpha}$ also given by Eq. (53).
The predictions for the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ depend on the parameter regime. For the case with $n>1$ and $(4n^{2}-2n\sqrt{4n^{2}-1})/3$ << $\tilde{\alpha}<{4n^{2}}/[3(n^{2}-1)]$, $1/\sqrt{3}<n<1$ and $\tilde{\alpha}>(4n^{2}-2n\sqrt{4n^{2}-1})/3$ or $1/3<n<1/\sqrt{3}$ and $\tilde{\alpha}>4n^{2}/[3(9n^{2}-1)]$, the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ are given by Kallosh:2013yoa ; Yi:2016jqr
$n_{s}=1-\frac{2}{N}+\frac{2NA-6n(N-1)}{N\left[2NA+n\left(4N^{2}/\tilde{\alpha}+3\right)\right]},$ (62)
$r=\frac{48(1-\lambda)n}{2NA+n\left(4N^{2}/\tilde{\alpha}+3\right)},$ (63)
where $A=\sqrt{12n^{2}/\tilde{\alpha}+9}$. For the case with $n>1$ and $\tilde{\alpha}>{4n^{2}}/[3(n^{2}-1)]$, the expressions become Kallosh:2013yoa ; Yi:2016jqr
$n_{s}=1-\frac{2}{N}+\frac{8n\left[(N-1)B+3\tilde{\alpha}n\right]}{N\left[\left(B+4nN+3\tilde{\alpha}n\right)^{2}-9\tilde{\alpha}^{2}\right]},$ (64)
$r=\frac{192(1-\lambda)\tilde{\alpha}n^{2}}{\left(B+4nN+3\tilde{\alpha}n\right)^{2}-9\tilde{\alpha}^{2}},$ (65)
where $B=\sqrt{9\tilde{\alpha}^{2}+24\tilde{\alpha}n^{2}+4n^{2}}-3\tilde{\alpha}n-2n$. For all other parameter choices, the predictions are given by Kallosh:2013yoa ; Yi:2016jqr
$n_{s}=1-\frac{2}{N}+\frac{8n\left[(N+1)C-3n\tilde{\alpha}\right]}{N\left[\left(C+4nN+3\tilde{\alpha}n\right)^{2}-9\tilde{\alpha}^{2}\right]},$ (66)
$r=\frac{192(1-\lambda)\tilde{\alpha}n^{2}}{\left(C+4nN+3\tilde{\alpha}n\right)^{2}-9\tilde{\alpha}^{2}},$ (67)
where $C=\sqrt{9\tilde{\alpha}^{2}-24\tilde{\alpha}n^{2}+4n^{2}}-3\tilde{\alpha}n+2n$.
Similar to the E-model, in the small-$\tilde{\alpha}$ limit, the predictions of the T-model reduce to those of the $\alpha$-attractor form given in Eq. (59). In the large-$\tilde{\alpha}$ limit, the T-model approaches the chaotic inflation model with a power-law potential of index $p=2n$ Kallosh:2013yoa , and the predictions coincide with those in Eq. (49). Taking $N=60$, the observational constraints on the parameters $\tilde{\alpha}$ and $n$ of the T-model from the P-ACT-LB data are shown in Fig. 4. For $n=1/2$, the predictions of the T-model for different values of $\tilde{\alpha}$ are shown in Fig. 2, represented by the red curves. With the inclusion of the Gauss-Bonnet term and a coupling constant $\lambda=0.8$, the T-model becomes consistent with the P-ACT-LB-BK18 observational data for sufficiently large $\tilde{\alpha}$.
Figure 4: Constraints on the parameters $n$ and $\tilde{\alpha}$ of the T-model, derived from the P-ACT-LB data as given in Eq. (46). The blue region corresponds to parameter values for which the predicted scalar spectral index given by Eq. (62) from the T-model is consistent with the P-ACT-LB data.